I'm trying to help my daughter and I'm stuck. I think the answer is x=-13.5 (i used a calculator) but I can't it this answer by hand. Help me with 2x-7/3=4/5
Answer by scott8148(6628)

2x=4/5+7/3 ... 2x=12/15+35/15 ... 2x=47/15 ... x=47/30